# CausalProfiler: Generating Synthetic Benchmarks for Rigorous and Transparent Evaluation of Causal Machine Learning

### v1

@@ 6.5 Experiment 2: Counterfactual Estimation on Discrete SCMs
This experiment evaluates counterfactual estimation on discrete-variable SCMs as a robustness check, testing CausalNF and DCM—originally designed for continuous settings—motivated by prior work showing that CausalNF can approximate discrete distributions (javaloy2023causal; manela2024marginal). We consider three discrete SoIs: Disc-C2-Reject, with 10-15 nodes, binary variables, and rejection-based mechanism sampling; Disc-C4-Unbias, with the same graph size but 4-category variables and unbiased random mechanism sampling; and Disc-L-C2-Unbias, with larger graphs (20-30 nodes), binary variables, and unbiased random mechanism sampling (Table 2).
@@ 7 Limitations and Future Work
Bridging the simulation-to-real gap. While synthetic evaluation is indispensable (audrey2025), it is insufficient to fully assess method capabilities, as results may not transfer to real-world settings. In CausalProfiler, alignment with real domains currently relies on manually specified SoIs, guided by domain expertise or empirical features. A key direction for future work is to develop methods that automatically map real data to sets of SoIs, enabling principled semi-synthetic evaluation pipelines where SoIs are shaped by empirical evidence rather than fixed assumptions. However, mapping from observational data to SoIs is a fundamentally underconstrained problem, and any such inference must be handled with care, given the challenges around identifiability and inductive bias.
@@ Appendix B Additional definitions & Notations
An SCM is said to be semi-Markovian (pearl09) if its set of structural equations is acyclic, meaning there exists an ordering of the equations such that for any two functions $f_{i},f_{j}\in\mathbfcal{F}$, if $f_{i}<f_{j}$, then $V_{j}\notin\textit{{PA}}(V_{i})$. This condition ensures that the causal dependencies among endogenous variables form a Directed Acyclic Graph.
An SCM is Markovian (pearl09) if the exogenous variables influencing different endogenous variables are mutually independent. Formally, for all distinct $V_{i},V_{j}\in\mathbf{V}$, we have $\mathbf{U}_{V_{i}}\perp\!\!\!\perp\mathbf{U}_{V_{j}}$. This implies the absence of latent confounding, allowing the model to be fully described by a DAG with independent noise terms.
@@
The causal graph of a Semi-Markovian (bareinboim:22) SCM is an acyclic directed mixed graph with:
@@ B.2 Counterfactual Quantities ($\mathcal{L}_{3}$)
A counterfactual query such as $P(Y_{\textit{{do}}(T=t)}|\mathbf{V}_{F}=\mathbf{v}_{F})$ is computed by abduction (conditioning on factual data), action (intervening), and prediction (computing the outcome) (pearl09).
@@
Originally, (zhang:18) defined counterfactual direct, indirect, and spurious effects by conditioning on the factual realization of one variable. Later, (plecko:24) generalized this to allow the factual evidence to be any subset $\mathbf{V}_{F}$ of endogenous variables, enabling more granular and flexible counterfactual analyses.
@@ C.2.1 Example 1: Price elasticity
The team wants to calculate discrete elasticity using non-parametric DoubleML methods (chernozhukov2018double). However, they do not know which method is more suitable for their setting. Hence, the team decides to use CausalProfiler to perform their own comparison and define the following set of SoI parameters:
@@ Appendix D Causal Graph Sampling
Algorithm 3 Projection Algorithm (verma93)
@@ E.1 Regional Discrete SCMs
The class of regional discrete SCMs got inspired by the class of Regional Canonical Models by (xia2023neural) and the class of canonical SCMs by (zhang22). We decided to define our own class rather than using one of these two classes for two reasons. First, canonical SCMs are very expensive to sample particularly because of the presence of confounded components. Second, even if Regional Canonical Models are designed to be less expensive because their expressivity can be regulated via the number of noise regions to consider, they lose some interesting properties such as the non overlapping of the noise regions which is crucial to favor a strong dependence between the user choice of the number of noise regions and the complexity of the generated mechanisms. Moreover, Regional Canonical Models still rely on confounded components which is the major source of complexity at the sampling stage. Hence, we defined the class of Regional discrete SCMs to not have to deal with confounded components at the sampling stage (instead we rely on a projection algorithm after sampling, see Appendix D) and to regulate mechanisms expressivity through the use of non-overlapping noise regions.
@@ Appendix G Analysis module’s metrics
  • Number of maximal confounded components (c-comps) [^5] We use (tian2002general) definition of (maximal) confounded components.: $|\mathbf{C}|$ where $\mathbf{C}$ denotes the set of maximal c-comps in $\mathcal{G}_{\mathbf{V}_{o}}$
@@ H.3 Comparision to CausalNF synthetic SCMs used for evaluation
To illustrate the contribution in SCMs diversity that CausalProfiler can give to practitioners wishing to evaluate Causal ML methods, we compare the SCMs sampled in the previous Section with those used in the CausalNF work (javaloy2023causal) for evaluation. We decided to first focus on the CausalNF synthetic SCMs because they have been reused by other papers (tram_dag; ccnf) to evaluate new methods as if they were classical synthetic benchmarks for counterfactual evaluation.
@@
For this comparison, we reimplemented the synthetic SCMs of CausalNF using CausalProfiler, and applied all the metrics of the analysis module (cf. Appendix G). In this way, the CausalNF SCMs were processed in the same way as our SCMs. We then used these metrics to compare the two groups of SCMs. For the sake of having a fair comparison, not penalizing the fact that some assumptions were taken by the authors, we removed some metrics from the analysis: the hidden confounders and positivity metrics. Indeed, all CausalNF SCMs satisfy the causal sufficiency and strong positivity hypotheses, whereas, as presented in Appendix H.2, our SCMs do not by design. Finally, in order to obtain an easily interpretable visual result, we applied a two-dimensional t-SNE projection (maaten2008visualizing) to all these metrics and subgroups of metrics (Figure 10). Each t-SNE has been applied here with a perplexity of 30.
@@ H.4 Comparision to bnlearn semi-sythetic graphical causal models
This section also illustrates the contribution CausalProfiler makes to SCMs’ diversity by comparing them with other causal models used in the literature: CANCER and EARTHQUAKE from bnlearn (bnlearn). Unlike the synthetic and continuous SCMs from CausalNF, CANCER and EARTHQUAKE are discrete causal graph models. The following analysis, therefore, enriches the conclusions of the previous section.
@@
Further, for this analysis we require datasets whose characteristics match those of the studied SoIs, in particular datasets with a small number of variables (3-5). This is why we include CausalNF and bnlearn networks, and exclude IHDP (hill2011bayesian), Twins (louizos2017causal), Syntren (syntren_06), and ACIC2016 (dorie2019automated), which contain substantially more variables. One might argue that we could simply sample higher-dimensional SCMs from CausalProfiler. While this is possible, computing the full set of assumption-analysis metrics (Appendix G) becomes computationally expensive as dimensionality and graph density increase; for example, Markov property checks and pairwise independence tests scale poorly with the number of variables. As a result, performing a detailed comparison with higher-dimensional datasets is not very tractable.
@@ Appendix I Visual overview of CausalProfiler’s sampling strategy
Remark on continuous SCMs. The universal approximation theorem (HORNIK1991251) states that NNs (with non-polynomial activation functions) are dense in the space of continuous functions, meaning that any continuous function can be approximated by a sequence of NNs converging to this function. However, this does not guarantee that they strictly cover the space of continuous functions. In particular, whenever the number of layers and neurons is finite, one can always build a continuous function too complex to be represented with this finite number of parameters. Hence, Proposition 5.1 cannot be extended to any class of continuous SCMs. However, it could potentially be adapted not to ask for strict coverage but rather density. We leave this question for future work.
@@
All following experiments are done on discrete SCMs to reduce approximations. Indeed, distributions over continuous variables can only be approximated (e.g., using kernel methods) while discrete ones can be computed exactly. In addition, the experiments rely on conditional independence testing, which has been proven to be particularly difficult to use with continuous variables. Indeed, (Shah_2020) proved that no conditional independence test with a continuous conditioning variable can have both a valid significance level and power.
@@
As the sampled SCMs are regional discrete, the conditional independence $A\perp\!\!\!\perp_{P_{\mathcal{M}}}B|\mathbf{C}$ can be tested with Pearson’s $\chi^{2}$ independence tests (pearson1900). More precisely, $A$ and $B$ are considered independent conditionally to $\mathbf{C}$ if for all values $\mathbf{c}$ of $\mathbf{C}$, the $H_{0}$ hypothesis "$A$ and $B$ are independent" is not rejected. Since Pearson’s $\chi^{2}$ test is based on the assumption that the number of samples is large, we decide to skip tests where the Koehler criterion (koehler_80) is not met. Based on empirical analyses, this criterion indicates whether the $\chi^{2}$ test is reliable depending on the number of samples considered. In addition, as we conduct tests for each observed value $c$, we need to control for the expected proportion of false positives (represented by the Type I error of the test). To do so, we apply the Benjamini-Hochberg correction (benjamini1995controlling).
@@
$|\mathbf{C}|=1$
$100$
(2 3912\,391)
$91.76$
(2 1942\,194)
$4.94$
($118$)
$3.3$
($79$)
$100$
(9 1309\,130)
$85.4$
(7 7977\,797)
$1.43$
($131$)
$13.17$
(1 2021\,202)
$|\mathbf{C}|=2$
$100$
(2 9862\,986)
$91.16$
(2 7222\,722)
$5.63$
($168$)
$3.22$
($96$)
$100$
(53 04053\,040)
$45.2$
(23 97623\,976)
$0.33$
($177$)
$54.46$
(28 88728\,887)
$|\mathbf{C}|=3$
$100$
(1 6931\,693)
$91.08$
(1 5421\,542)
$5.67$
($96$)
$3.25$
($55$)
$100$
(145 320145\,320)
$18.49$
(26 87426\,874)
$0.07$
($106$)
$81.43$
(118 340118\,340)
Total
$100$
(7 0707\,070)
$91.34$
(6 4586\,458)
$5.40$
($382$)
$3.25$
($230$)
$100$
(207 490207\,490)
$28.26$
(58 64758\,647)
$0.2$
($414$)
$71.54$
(148 429148\,429)
@@
Do-Calculus rules (pearl09)
@@
For the same computational cost reasons as for $\mathcal{L}_{1}$ verification, we consider only univariate sets of variables $X,Y,Z$, and $W$. In addition, the studied SCMs are sampled from the same SoIs as defined in the $\mathcal{L}_{1}$-verification experiment (Appendix K.1.1). Finally, to assess whether two conditional distributions are identical, we used Pearson’s $\chi^{2}$ goodness of fit tests (pearson1900). As done in Section K.1, we also use the Koehler criterion (koehler_80) and the Benjamini-Hochberg correction (benjamini1995controlling).
@@
Rule 1
Insertion/deletion
of observation
$100$
(3 3783\,378)
$96.15$
(3 2483\,248)
$3.85$
($130$)
0
$100$
(171 092171\,092)
$88.84$
(152 004152\,004)
$0.1$
($172$)
$11.06$
(18 91618\,916)
Rule 2
Action/observation
exchange
$100$
(5 0655\,065)
$94.04$
(4 7634\,763)
$5.96$
($302$)
0
$100$
(259 509259\,509)
$83.84$
(217 578217\,578)
$0.09$
($241$)
$16.06$
(41 69041\,690)
Rule 3
Insertion/deletion
of action
$100$
(5 1695\,169)
$93.75$
(4 8464\,846)
$6.25$
($323$)
0
$100$
(282 184282\,184)
$89.21$
(251 731251\,731)
$0.06$
($157$)
$10.74$
(30 29630\,296)
Total
$100$
(13 61213\,612)
$94.45$
(12 85712\,857)
$5.55$
($755$)
0
$100$
(712 785712\,785)
$87.17$
(621 313621\,313)
$0.08$
($570$)
$12.75$
(90 90290\,902)
@@
Axiomatic characterization of structural counterfactuals (pearl09)
@@
We evaluate three methods: CausalNF (javaloy2023causal), DCM (chao2023interventional), and DeCaFlow (almodovar2025decaflow). The first two methods assume causal sufficiency, and therefore cannot, in theory, handle hidden confounding. DeCaFlow, in contrast, is explicitly designed for this setting but requires access to the full causal graph (including hidden variables) and does not run when all variables are observed. Thus, we include it only in the hidden confounding SoI.

### v2

@@ 6.5 Experiment 2: Counterfactual Estimation on Discrete SCMs
This experiment evaluates counterfactual estimation on discrete-variable SCMs as a robustness check, testing CausalNF and DCM—originally designed for continuous settings—motivated by prior work showing that CausalNF can approximate discrete distributions (Javaloy et al., 2023; de Vassimon Manela et al., 2024). We consider three discrete SoIs: Disc-C2-Reject, with 10-15 nodes, binary variables, and rejection-based mechanism sampling; Disc-C4-Unbias, with the same graph size but 4-category variables and unbiased random mechanism sampling; and Disc-L-C2-Unbias, with larger graphs (20-30 nodes), binary variables, and unbiased random mechanism sampling (Table 2).
@@ 7 Limitations and Future Work
Bridging the simulation-to-real gap. While synthetic evaluation is indispensable (Poinsot et al., 2025), it is insufficient to fully assess method capabilities, as results may not transfer to real-world settings. In CausalProfiler, alignment with real domains currently relies on manually specified SoIs, guided by domain expertise or empirical features. A key direction for future work is to develop methods that automatically map real data to sets of SoIs, enabling principled semi-synthetic evaluation pipelines where SoIs are shaped by empirical evidence rather than fixed assumptions. However, mapping from observational data to SoIs is a fundamentally underconstrained problem, and any such inference must be handled with care, given the challenges around identifiability and inductive bias.
@@ 8 Conclusion
Acknowledgements
This research was partially funded by the French National Research Agency (ANR) under the France 2030 program, under the reference 23-PEIA-004, by the Artificial Intelligence for Safe and Smart Mobility Chair (Grant No. ANR-23-CPJ1-0099-01), and by the UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI (ART-AI) [EP/S023437/1].
@@ Appendix B Additional definitions & Notations
An SCM is said to be semi-Markovian (Pearl, 2009) if its set of structural equations is acyclic, meaning there exists an ordering of the equations such that for any two functions $f_{i},f_{j}\in\mathbfcal{F}$, if $f_{i}<f_{j}$, then $V_{j}\notin\textit{{PA}}(V_{i})$. This condition ensures that the causal dependencies among endogenous variables form a Directed Acyclic Graph.
An SCM is Markovian (Pearl, 2009) if the exogenous variables influencing different endogenous variables are mutually independent. Formally, for all distinct $V_{i},V_{j}\in\mathbf{V}$, we have $\mathbf{U}_{V_{i}}\perp\!\!\!\perp\mathbf{U}_{V_{j}}$. This implies the absence of latent confounding, allowing the model to be fully described by a DAG with independent noise terms.
@@
The causal graph of a Semi-Markovian (Bareinboim et al., 2022) SCM is an acyclic directed mixed graph with:
@@ B.2 Counterfactual Quantities ($\mathcal{L}_{3}$)
A counterfactual query such as $P(Y_{\textit{{do}}(T=t)}|\mathbf{V}_{F}=\mathbf{v}_{F})$ is computed by abduction (conditioning on factual data), action (intervening), and prediction (computing the outcome) (Pearl, 2009).
@@
Originally, (Zhang and Bareinboim, 2018) defined counterfactual direct, indirect, and spurious effects by conditioning on the factual realization of one variable. Later, (Plečko and Bareinboim, 2024) generalized this to allow the factual evidence to be any subset $\mathbf{V}_{F}$ of endogenous variables, enabling more granular and flexible counterfactual analyses.
@@ C.2.1 Example 1: Price elasticity
The team wants to calculate discrete elasticity using non-parametric DoubleML methods (Chernozhukov et al., 2018). However, they do not know which method is more suitable for their setting. Hence, the team decides to use CausalProfiler to perform their own comparison and define the following set of SoI parameters:
@@ Appendix D Causal Graph Sampling
Algorithm 3 Projection Algorithm (Verma, 1993)
@@ E.1 Regional Discrete SCMs
The class of regional discrete SCMs got inspired by the class of Regional Canonical Models by (Xia et al., 2023) and the class of canonical SCMs by (Zhang et al., 2022). We decided to define our own class rather than using one of these two classes for two reasons. First, canonical SCMs are very expensive to sample particularly because of the presence of confounded components. Second, even if Regional Canonical Models are designed to be less expensive because their expressivity can be regulated via the number of noise regions to consider, they lose some interesting properties such as the non overlapping of the noise regions which is crucial to favor a strong dependence between the user choice of the number of noise regions and the complexity of the generated mechanisms. Moreover, Regional Canonical Models still rely on confounded components which is the major source of complexity at the sampling stage. Hence, we defined the class of Regional discrete SCMs to not have to deal with confounded components at the sampling stage (instead we rely on a projection algorithm after sampling, see Appendix D) and to regulate mechanisms expressivity through the use of non-overlapping noise regions.
@@ Appendix G Analysis module’s metrics
  • Number of maximal confounded components (c-comps) [^5] We use (Tian and Pearl, 2002) definition of (maximal) confounded components.: $|\mathbf{C}|$ where $\mathbf{C}$ denotes the set of maximal c-comps in $\mathcal{G}_{\mathbf{V}_{o}}$
@@ H.3 Comparision to CausalNF synthetic SCMs used for evaluation
To illustrate the contribution in SCMs diversity that CausalProfiler can give to practitioners wishing to evaluate Causal ML methods, we compare the SCMs sampled in the previous Section with those used in the CausalNF work (Javaloy et al., 2023) for evaluation. We decided to first focus on the CausalNF synthetic SCMs because they have been reused by other papers (Sick and Dürr, 2025; Zhou et al., 2025) to evaluate new methods as if they were classical synthetic benchmarks for counterfactual evaluation.
@@
For this comparison, we reimplemented the synthetic SCMs of CausalNF using CausalProfiler, and applied all the metrics of the analysis module (cf. Appendix G). In this way, the CausalNF SCMs were processed in the same way as our SCMs. We then used these metrics to compare the two groups of SCMs. For the sake of having a fair comparison, not penalizing the fact that some assumptions were taken by the authors, we removed some metrics from the analysis: the hidden confounders and positivity metrics. Indeed, all CausalNF SCMs satisfy the causal sufficiency and strong positivity hypotheses, whereas, as presented in Appendix H.2, our SCMs do not by design. Finally, in order to obtain an easily interpretable visual result, we applied a two-dimensional t-SNE projection (Maaten and Hinton, 2008) to all these metrics and subgroups of metrics (Figure 10). Each t-SNE has been applied here with a perplexity of 30.
@@ H.4 Comparision to bnlearn semi-sythetic graphical causal models
This section also illustrates the contribution CausalProfiler makes to SCMs’ diversity by comparing them with other causal models used in the literature: CANCER and EARTHQUAKE from bnlearn (Scutari, 2019). Unlike the synthetic and continuous SCMs from CausalNF, CANCER and EARTHQUAKE are discrete causal graph models. The following analysis, therefore, enriches the conclusions of the previous section.
@@
Further, for this analysis we require datasets whose characteristics match those of the studied SoIs, in particular datasets with a small number of variables (3-5). This is why we include CausalNF and bnlearn networks, and exclude IHDP (Hill, 2011), Twins (Louizos et al., 2017), Syntren (den Bulcke et al., 2006), and ACIC2016 (Dorie et al., 2019), which contain substantially more variables. One might argue that we could simply sample higher-dimensional SCMs from CausalProfiler. While this is possible, computing the full set of assumption-analysis metrics (Appendix G) becomes computationally expensive as dimensionality and graph density increase; for example, Markov property checks and pairwise independence tests scale poorly with the number of variables. As a result, performing a detailed comparison with higher-dimensional datasets is not very tractable.
@@ Appendix I Visual overview of CausalProfiler’s sampling strategy
Remark on continuous SCMs. The universal approximation theorem (Hornik, 1991) states that NNs (with non-polynomial activation functions) are dense in the space of continuous functions, meaning that any continuous function can be approximated by a sequence of NNs converging to this function. However, this does not guarantee that they strictly cover the space of continuous functions. In particular, whenever the number of layers and neurons is finite, one can always build a continuous function too complex to be represented with this finite number of parameters. Hence, Proposition 5.1 cannot be extended to any class of continuous SCMs. However, it could potentially be adapted not to ask for strict coverage but rather density. We leave this question for future work.
@@
All following experiments are done on discrete SCMs to reduce approximations. Indeed, distributions over continuous variables can only be approximated (e.g., using kernel methods) while discrete ones can be computed exactly. In addition, the experiments rely on conditional independence testing, which has been proven to be particularly difficult to use with continuous variables. Indeed, (Shah and Peters, 2020) proved that no conditional independence test with a continuous conditioning variable can have both a valid significance level and power.
@@
As the sampled SCMs are regional discrete, the conditional independence $A\perp\!\!\!\perp_{P_{\mathcal{M}}}B|\mathbf{C}$ can be tested with Pearson’s $\chi^{2}$ independence tests (Pearson, 1900). More precisely, $A$ and $B$ are considered independent conditionally to $\mathbf{C}$ if for all values $\mathbf{c}$ of $\mathbf{C}$, the $H_{0}$ hypothesis "$A$ and $B$ are independent" is not rejected. Since Pearson’s $\chi^{2}$ test is based on the assumption that the number of samples is large, we decide to skip tests where the Koehler criterion (Koehler and Larntz, 1980) is not met. Based on empirical analyses, this criterion indicates whether the $\chi^{2}$ test is reliable depending on the number of samples considered. In addition, as we conduct tests for each observed value $c$, we need to control for the expected proportion of false positives (represented by the Type I error of the test). To do so, we apply the Benjamini-Hochberg correction (Benjamini and Hochberg, 1995).
@@
$|\mathbf{C}|=1$ $100$ (2 3912\,391) $91.76$ (2 1942\,194) $4.94$ ($118$) $3.3$ ($79$) $100$ (9 1309\,130) $85.4$ (7 7977\,797) $1.43$ ($131$) $13.17$ (1 2021\,202)
$|\mathbf{C}|=2$ $100$ (2 9862\,986) $91.16$ (2 7222\,722) $5.63$ ($168$) $3.22$ ($96$) $100$ (53 04053\,040) $45.2$ (23 97623\,976) $0.33$ ($177$) $54.46$ (28 88728\,887)
$|\mathbf{C}|=3$ $100$ (1 6931\,693) $91.08$ (1 5421\,542) $5.67$ ($96$) $3.25$ ($55$) $100$ (145 320145\,320) $18.49$ (26 87426\,874) $0.07$ ($106$) $81.43$ (118 340118\,340)
Total $100$ (7 0707\,070) $91.34$ (6 4586\,458) $5.40$ ($382$) $3.25$ ($230$) $100$ (207 490207\,490) $28.26$ (58 64758\,647) $0.2$ ($414$) $71.54$ (148 429148\,429)
@@
Do-Calculus rules (Pearl, 2009)
@@
For the same computational cost reasons as for $\mathcal{L}_{1}$ verification, we consider only univariate sets of variables $X,Y,Z$, and $W$. In addition, the studied SCMs are sampled from the same SoIs as defined in the $\mathcal{L}_{1}$-verification experiment (Appendix K.1.1). Finally, to assess whether two conditional distributions are identical, we used Pearson’s $\chi^{2}$ goodness of fit tests (Pearson, 1900). As done in Section K.1, we also use the Koehler criterion (Koehler and Larntz, 1980) and the Benjamini-Hochberg correction (Benjamini and Hochberg, 1995).
@@
Rule 1 Insertion/deletion of observation $100$ (3 3783\,378) $96.15$ (3 2483\,248) $3.85$ ($130$) 0 (0) $100$ (171 092171\,092) $88.84$ (152 004152\,004) $0.1$ ($172$) $11.06$ (18 91618\,916)
Rule 2 Action/observation exchange $100$ (5 0655\,065) $94.04$ (4 7634\,763) $5.96$ ($302$) 0 (0) $100$ (259 509259\,509) $83.84$ (217 578217\,578) $0.09$ ($241$) $16.06$ (41 69041\,690)
Rule 3 Insertion/deletion of action $100$ (5 1695\,169) $93.75$ (4 8464\,846) $6.25$ ($323$) 0 (0) $100$ (282 184282\,184) $89.21$ (251 731251\,731) $0.06$ ($157$) $10.74$ (30 29630\,296)
Total $100$ (13 61213\,612) $94.45$ (12 85712\,857) $5.55$ ($755$) 0 (0) $100$ (712 785712\,785) $87.17$ (621 313621\,313) $0.08$ ($570$) $12.75$ (90 90290\,902)
@@
Axiomatic characterization of structural counterfactuals (Pearl, 2009)
@@
We evaluate three methods: CausalNF (Javaloy et al., 2023), DCM (Chao et al., 2023), and DeCaFlow (Almodóvar et al., 2025). The first two methods assume causal sufficiency, and therefore cannot, in theory, handle hidden confounding. DeCaFlow, in contrast, is explicitly designed for this setting but requires access to the full causal graph (including hidden variables) and does not run when all variables are observed. Thus, we include it only in the hidden confounding SoI.